OpenStudy (anonymous):

The area of a triangle is 48 square inches, and its height is 16 inches. Find the base of the triangle.

OpenStudy (solomonzelman):

the formula is A= (b × h) / 2 so substitute the numbers and solve! 48= (b × 16) /2 48= 16b /2 = 8b the answer is 6=b (medal plizzz?)

OpenStudy (solomonzelman):

Need more explanations? I'll explain!

OpenStudy (anonymous):

thanks ! :)